OpenStudy (anonymous):

Column #1 The x-coordinate of the vertex of the equation y = 2x2 − 4x + 12 Column #2 The x-coordinate of the vertex of the equation y = 4x2 + 8x + 3 a. The value found in column #1 is greater than the value found in column #2. b. The value found in column #1 is less than the value found in column #2. c. The value found in column #1 is equivalent to the value found in column #2. d. The relationship between column #1 and column #2 cannot be determined by the information given.

TheSmartOne (thesmartone):

The x-coordinate of the vertex for the equation: \(\sf \color{bl}{a}x^2+\color{red}{b}x+c\) is \(\sf\Large h=\frac{-\color{red}{b}}{2\color{blue}{a}}\) The h is because the vertex is written as \(\sf (h,k)\) so the x-coordinate of the vertex would be 'h'.

OpenStudy (anonymous):

Thank you so much!

TheSmartOne (thesmartone):

The x-coordinate of the vertex for the equation: \(\sf \color{blue}{a}x^2+\color{red}{b}x+c\) is \(\sf\Large h=\frac{-\color{red}{b}}{2\color{blue}{a}}\) The h is because the vertex is written as \(\sf (h,k)\) so the x-coordinate of the vertex would be 'h'. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ I fixed it up to include the blue 'a' in ax^2+bx+c as I messed up with the latex :p